OpenStudy (anonymous):

Find the length of the missing side. b = 2.1, c = 2.9

OpenStudy (anonymous):

yeah

OpenStudy (jiteshmeghwal9):

a+b+c=180 (angle sum property) a+2.1+2.9=180 a+5=180 solve it for 'a' :)